OpenStudy (anonymous):

can someone help me here it is two-fifths times seven-ninths equals what fraction?

OpenStudy (anonymous):

Multiply both numerators together and both denominators together to get your answer

OpenStudy (anonymous):

i have no idea how to do that

OpenStudy (anonymous):

\[\frac{ 2 }{ 5 } \times \frac{ 7 }{ 9 } = \frac{ 2 \times 7 }{ 5 \times 9 }\]

OpenStudy (anonymous):

oh ty so much

OpenStudy (anonymous):

No problem